# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ 7.1 Construction of the GMM Approximation of a SNN
$\bar{{\mathcal{C}}}(N)$ $=\underset{{\mathcal{C}}\in{\mathbb{R}},|{\mathcal{C}}|=N}{\mathrm{argmin}}{\mathbb{W}}_{2}({\Delta}_{{\mathcal{C}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))$ (26)
$\{N_{1}^{*},\ldots,N_{n}^{*}\}$ $=\underset{\begin{subarray}{c}\{N_{1},\ldots,N_{n}\}\in\mathbb{N}^{n},\\ \prod_{j=1}^{n},N_{j}=N\end{subarray}}{\mathrm{argmin}}\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}_{2}({\Delta}_{\bar{{\mathcal{C}}}(N_{i})}\#\mathcal{N}(0,1),\mathcal{N}(0,1)).$ (27)
Following Proposition 19, the optimal grid of signature locations of any Gaussian is defined by the optimal signature locations for a univariate Gaussian (Eqn. (26)) and the optimal size of each dimension of the grid (Eqn. (27)). While the optimization problem in Eqn. (26) is non-convex, it can be solved up to any precision using the fixed-point approach proposed in Kieffer (1982). In particular, we solve the problem once for a range of grid sizes $N$ and store the optimal locations in a lookup table that is called at runtime. To address the optimization problem in Eqn. (27), we use that ${\mathbb{W}}_{2}({\Delta}_{\bar{{\mathcal{C}}}(N)}\#\mathcal{N}(0,1),\mathcal{N}(0,1))$, with optimal locations $\bar{{\mathcal{C}}}(N)$ as in Eqn. (26), is strictly decreasing for increasing $N$ so that even for large $N$, the number of feasible non-dominated candidates $\{N_{1},\ldots,N_{n}\}$ is small.
input : $\sum_{i=1}^{M}{\bar{{{\pi}}}^{{(i)}}}\mathcal{N}({m}_{i},\Sigma_{i})$
output : $d=\sum_{j=1}^{N}{{{\pi}}^{(j)}}\delta_{{{c}}_{j}}$; $\hat{{\mathbb{W}}}_{2,{\Delta}}$
@@ 1 Compute eigenvalues vector ${{\lambda}}_{i}$ and eigenvector matrix ${{V}}_{i}$ of $\Sigma_{i}$
            2 Find the optimal grid sizes $\{N_{1}^{*},\ldots,N_{n}^{*}\}$ for ${{\lambda}}_{i}$ according to Eqn. (26)
            3 Take grid $\bar{{\mathcal{C}}}=\bar{{\mathcal{C}}}(N_{1})\times\ldots\times\bar{{\mathcal{C}}}(N_{n})$ , with $\bar{{\mathcal{C}}}(N_{i})$ as in Eqn. (27)
            4 Transform the grid to the original space: ${\mathcal{C}}_{i}=\mathrm{Post}(\bar{{\mathcal{C}}},{{S}})+{m}_{i}$ with ${{S}}={{V}}_{i}\mathrm{diag}({{\lambda}}_{i})^{\frac{1}{2}}$
            5 Define $d_{i}=\sum_{j=1}^{|{\mathcal{C}}_{i}|}{{{\pi}}^{{(j)}}}{{c}}_{j}$ where ${{c}}_{j}\in{\mathcal{C}}_{i}$ and ${{{\pi}}^{{(j)}}}=\mathbb{P}_{{{z}}\sim\mathcal{N}({m}_{i},\Sigma_{i})}[{{z}}\in{\mathcal{R}}_{j}]$
            6 Compute $\hat{{\mathbb{W}}}_{2,{\Delta},i}=\sum_{j=1}^{|{\mathcal{C}}|}{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{j}}{{{\pi}}^{(j)}}\hat{{\mathbb{W}}}_{2\mid{\mathcal{R}}_{i}}$ with $\hat{{\mathbb{W}}}_{2\mid{\mathcal{R}}_{j}}=\sum_{l=1}^{n}{{{\lambda}}_{i}^{{(l)}}}({{{\nu}}_{j}^{{(l)}}}+({{{\mu}}_{j}^{{(l)}}}-{{{c}}_{j}^{{(l)}}}))$ with ${{{\mu}}_{j}^{{(l)}}}$ as in Eqn. (8) and ${{{\nu}}_{j}^{{(l)}}}$ as in Eqn. (9) of Prop. 10
      7Set $d=\sum_{i=1}^{M}{\bar{{{\pi}}}^{{(i)}}}d_{i}$ and $\hat{{\mathbb{W}}}_{2,{\Delta}}^{2}=\sum_{i=1}^{M}{\bar{{{\pi}}}^{{(i)}}}{\mathbb{W}}_{2,{\Delta},i}^{2}$
@@
Compression Operation
Our approach to compress a GMM of size $N$ into a GMM of size $M<N$, i.e., operation ${R}_{M}$ in Eqn. (13b), is described in Algorithm 3 and is based on the M-means clustering using Lloyd’s algorithm (Lloyd, 1982) on the means of the mixture’s components (line 1). That is, each cluster is substituted by a Gaussian distribution with mean and covariance equal to those of the mixture in the cluster (line 2).
@@
For the compression of discrete distributions, such as in the case of Bayesian inference via dropout, we present a computationally more efficient alternative procedure to Algorithm 3 in Section B of the Appendix. Note that in the dropout case, the support of the multivariate Bernoulli distribution representing the dropout masks at each layer’s input grows exponentially with the layer width, making compression of paramount importance in practice.
input : $\bar{q}=\sum_{i=1}^{\bar{M}}{\bar{{{\pi}}}^{{(i)}}}\mathcal{N}(\bar{{m}}_{i},\bar{\Sigma}_{i})$
@@
      1 Collect the components of $\bar{q}$ in $M$ clusters by applying $M$-means clustering to $\{\bar{{m}}_{1},\ldots,\bar{{m}}_{\bar{M}}\}$ using Lloyd’s Algorithm
@@
            2 ${{{\pi}}^{(j)}}=\sum_{i\in{\mathcal{I}}_{j}}{\bar{{{\pi}}}^{{(i)}}}$; ${m}_{j}=\frac{1}{{{{\pi}}^{(j)}}}\sum_{i\in{\mathcal{I}}_{j}}{\bar{{{\pi}}}^{{(i)}}}\bar{{m}}_{i}$; $\Sigma_{j}=\frac{1}{{{{\pi}}^{(j)}}}\sum_{i\in{\mathcal{I}}_{j}}{\bar{{{\pi}}}^{{(i)}}}[\bar{\Sigma}_{i}+(\bar{{m}}_{i}-{m}_{j})(\bar{{m}}_{i}-{m}_{j})^{T}]$
@@ 7.2 Construction of a SNN Approximation of a GMM
The algorithmic framework presented in the previous subsection finds the Gaussian Mixture distribution that best approximates the output distribution of a SNN and returns bounds on their distance. In this subsection, we show how our results can be employed to solve the complementary problem of finding the weight distribution of a SNN to best match a given GMM.
Let us consider a finite set of input points ${\mathcal{X}}\subset{\mathbb{R}}^{n_{0}}$ and a SNN and GMM, whose distributions at ${\mathcal{X}}$ are respectively $p_{nn(\mathcal{X})}$ and $q_{gmm({\mathcal{X}})}$. Furthermore, in this subsection, we assume that the weight distribution $p_{{w}}$ of the SNN is parametrized by a set of parameters $\psi$. For instance, in the case where $p_{{w}}$ is a multivariate Gaussian distribution, then $\psi$ parametrizes its mean and covariance. Note that, consequently, also $p_{nn({\mathcal{X}})}$ depends on $\psi$ through Eqn. (21). To make the dependence explicit, in what follows, we will use a superscript to emphasize the dependency on $\psi$. Then, our goal is to find $\psi^{*}=\mathrm{argmin}_{\psi}{\mathbb{W}}_{2}(p_{nn(\mathcal{X})}^{\psi},q_{gmm(\mathcal{X})})$. To do so, we rely on Corollary 21 below, which uses the triangle inequality to extend Proposition 16.
Corollary 21 (of Proposition 16)
@@
${\mathbb{W}}_{2}(p_{nn({\mathcal{X}})}^{\psi},q_{gmm({\mathcal{X}})})\leq\hat{{\mathbb{W}}}_{2}(p_{nn({\mathcal{X}})}^{\psi},q_{nn({\mathcal{X}})}^{\psi})+{\mathbb{MW}}_{2}(q_{nn({\mathcal{X}})}^{\psi},q_{gmm({\mathcal{X}})}),$ (28)
where the GMM $q_{nn({\mathcal{X}})}^{\psi}$ and the bound $\hat{{\mathbb{W}}}_{2}$ are obtained via Algorithm 1, and ${\mathbb{MW}}_{2}$ is as defined in Definition 15, with $q_{nn({\mathcal{X}})}^{\psi}$ also depending on $\psi$ through Eqn. (22).
@@
[…]
where $\beta\in[0,1]$ allows us to trade off between the gap between the bound $\hat{{\mathbb{W}}}_{2}(p^{\psi}_{nn({\mathcal{X}})},q^{\psi}_{nn({\mathcal{X}})})$ and ${\mathbb{W}}_{2}(p^{\psi}_{nn({\mathcal{X}})},q^{\psi}_{nn({\mathcal{X}})})$ and the gap between the bound ${\mathbb{MW}}_{2}(q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})$ and ${\mathbb{W}}_{2}(q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})$. For instance, in the case where $q^{\psi}_{nn({\mathcal{X}})}$ and $q_{gmm({\mathcal{X}})}$ are Gaussian distributions, ${\mathbb{MW}}_{2}(q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})$ equals ${\mathbb{W}}_{2}(q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})$, leading us to choose $\beta<\frac{1}{2}$. As discussed next, under the assumption that the mean and covariance of $p_{{w}}$ are differentiable with respect to $\psi$, the objective in Eqn. (29) is piecewise differentiable with respect to $\psi$. Consequently, the optimization problem can be approximately solved using gradient-based optimization techniques, such as Adam (Kingma and Ba, 2014).
Let us first consider the term $\hat{{\mathbb{W}}}_{2}$ in the objective, which is iteratively defined in Eqn. (14) over the layers of the network using the quantities $\hat{\mathcal{S}}_{2,{{w}},k}$, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$, and $\hat{{\mathbb{W}}}_{2,{R},k}$ for each layer $k$. In Algorithm 1, we compute these quantities so that the gradients with respect to the mean and covariance of $p_{{w}}$ are (approximately) analytically tractable, as shown next:
@@
    $\hat{\mathcal{S}}_{2,{{w}},k}$ is computed according to Lemma 14 (line 6 of Algorithm 1) as the sum of the variances of the $k$-th layer’s weights and the spectral norm of the mean of the $k$-th layer’s weight matrix. Hence, the gradient with respect to the variance is well defined, and the gradient with respect to the mean exists if the largest singular value is unique; otherwise, we can use a sub-gradient derived from the singular value decomposition (SVD) of the mean matrix (Rockafellar, 2015).
@@
    $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ is obtained according to Algorithm 2 (line 4 of Algorithm 1) based on the eigendecomposition of the covariance matrix for each component of the GMM ${R}_{M}\#q_{nn({\mathcal{X}}),k}$ (line 1 of Algorithm 2). Differentiability of $\hat{{\mathbb{W}}}_{2,{\Delta},k}$ with respect to the mean and covariance of $p_{{w}}$ follows from the piecewise differentiability of the eigenvalue decomposition of a (semi-)positive definite matrix (Magnus and Neudecker, 2019), [^12] Note that there are discontinuous changes in the eigenvalues for small perturbations if the covariance matrix is (close to) degenerate, which occurs in our cases if the behavior of the SNN is strongly correlated for points in ${\mathcal{X}}$. To prevent this, we use a stable derivation implementation for degenerate covariance matrices as provided by Kasim (2020) to compute the eigenvalue decomposition. provided that the covariances of ${R}_{M}\#q_{nn({\mathcal{X}}),k}$ are differentiable with respect to the mean and covariance of $p_{{w}}$. To establish the latter, note that the compression operation ${R}_{M}$, as performed according to Algorithm 3, given the clustering of the components of the mixture $q_{nn({\mathcal{X}}),k}$, reduces to computing the weighted sum of the means and covariances of each cluster. Hence, the covariances of ${R}_{M}\#q_{nn({\mathcal{X}}),k}$ are piecewise differentiable with respect to the means and covariances of $q_{nn({\mathcal{X}}),k}$. Here, $q_{nn({\mathcal{X}}),k}$ is a linear combination of $p_{{{w}}_{k-1}}$ and ${{x}}$ if $k=1$, and the support of the signature approximation of ${R}_{M}\#q_{nn({\mathcal{X}}),k-1}$ otherwise (see Eqn. (13c) for the closed-form expression of $q_{nn({\mathcal{X}}),k}$ in the case ${\mathcal{X}}=\{{{x}}\}$). As such, from the chain rule, it follows that for all $k$ the means and covariances of ${R}_{M}\#q_{nn({\mathcal{X}}),k}$, and consequently, $\hat{{\mathbb{W}}}_{2,{\Delta},k}$, are piecewise differentiable with respect to the means and covariances of $p_{{{w}}_{k-1}},\ldots,p_{{{w}}_{0}}$.

### v2

@@ 7.1 Construction of the GMM Approximation of a SNN
${\mathcal{C}}^{*}(N)$ $=\underset{{\mathcal{C}}\in{\mathbb{R}},|{\mathcal{C}}|=N}{\mathrm{argmin}}{\mathbb{W}}_{2}\left({\Delta}_{{\mathcal{C}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right)$ (30)
$\{N_{1}^{*},\ldots,N_{n}^{*}\}$ $=\underset{\begin{subarray}{c}\{N_{1},\ldots,N_{n}\}\in\mathbb{N}^{n},\\ \prod_{j=1}^{n},N_{j}=N\end{subarray}}{\mathrm{argmin}}\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}_{2}\left({\Delta}_{{\mathcal{C}}^{*}(N_{i})}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right).$ (31)
Following Proposition 23, the optimal grid of signature locations of any Gaussian is defined by the optimal signature locations for a univariate Gaussian (Eqn (30)) and the optimal size of each dimension of the grid (Eqn (31)). While the optimization problem in Eqn (30) is non-convex, it can be solved up to any precision using the fixed-point approach proposed in Kieffer (1982). In particular, we solve the problem once for a range of grid sizes NN and store the optimal locations in a lookup table that is called at runtime. To address the optimization problem in Eqn (31), we use that ${\mathbb{W}}_{2}\left({\Delta}_{{\mathcal{C}}^{*}(N)}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right)$, with optimal locations ${\mathcal{C}}^{*}(N)$ as in Eqn (30), is strictly decreasing for increasing $N$ so that even for large $N$, the number of feasible non-dominated candidates $\{N_{1},\ldots,N_{n}\}$ is small.
input : ${q=}\sum_{i=1}^{M}{\tilde{{{\pi}}}^{{(i)}}}\mathcal{N}({m}_{i},\Sigma_{i})$
output : $d=\sum_{j=1}^{N}{{{\pi}}^{(j)}}\delta_{{{c}}_{j}}$ and $\hat{{\mathbb{W}}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}(q)$
@@ 1 Compute eigenvalues vector ${{\lambda}}_{i}$ and eigenvector matrix ${{V}}_{i}$ of $\Sigma_{i}$
    2 Find the optimal grid sizes $\{N_{1}^{*},\ldots,N_{n}^{*}\}$ for ${{\lambda}}_{i}$ according to Eqn (30)
    3 Construct grid of points ${\mathcal{C}}^{*}={\mathcal{C}}^{*}(N_{1})\times\ldots\times{\mathcal{C}}^{*}(N_{n})$ , with ${\mathcal{C}}^{*}(N_{i})$ as in Eqn (31)
    4 Transform the grid to the original space: ${\mathcal{C}}_{i}=\mathrm{Post}({\mathcal{C}}^{*},{{S}})+{m}_{i}$ where ${{S}}={{V}}_{i}\mathrm{diag}({{\lambda}}_{i})^{\frac{1}{2}}$
    5 Set $\hat{\mathbb{W}}_{2,{\Delta}_{{\mathcal{C}}_{i}}}^{2}=\sum_{l=1}^{n}{{{\lambda}}_{i}^{{(l)}}}\sum_{j=1}^{N_{l}}{{{\pi}}_{l}^{{(j)}}}\big{[}{{{\nu}}_{j}^{{(l)}}}+({{{\mu}}_{j}^{{(l)}}}-{\tilde{{c}}_{j}^{{(l)}}})^{2}\big{]}$, with ${\mathcal{C}}^{*}(N_{l})=\{{{{c}}_{j}^{{(l)}}}\}_{j=1}^{N_{l}}$, and ${{\pi}}_{l}^{{(j)}}$}, ${{\mu}}_{j}^{{(l)}}$}, and ${{\nu}}_{j}^{{(l)}}$} as in Eqn (11)
    6 Define $d_{i}=\sum_{j=1}^{|{\mathcal{C}}_{i}|}{{{\pi}}^{{(j)}}}{{c}}_{j}$ for ${{{\pi}}^{{(j)}}}=\prod_{l=1}^{n}{{{\pi}}_{l}^{{(j)}}}$
 7Return $d=\sum_{i=1}^{M}{\tilde{{{\pi}}}^{{(i)}}}d_{i}$ and $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}^{2}(q)=\sum_{i=1}^{M}{\tilde{{{\pi}}}^{{(i)}}}\hat{\mathbb{W}}_{2,{\Delta}_{{\mathcal{C}}_{i}}}^{2}$
@@
Algorithm 2 constructs a set of grid-constrained signatures that satisfy Eqn (12), thereby enabling formal error evaluation via Corollary 11. In settings where formal error bounds are not required, more flexible choices for the signature locations can alternatively be employed, provided they satisfy the conditions in Remark 7 that ensure tractability of the signature operations. One such alternative signature placement strategy is presented in Appendix D. In Section 8.1, we empirically compare the strategy in Algorithm 2 with the strategy in Appendix  D by estimating the Wasserstein distance using Monte Carlo methods.
input : $\tilde{q}=\sum_{i=1}^{\tilde{M}}{\tilde{{{\pi}}}^{{(i)}}}\mathcal{N}(\tilde{{m}}_{i},\tilde{\Sigma}_{i})$
@@
 1 Collect the components of $\tilde{q}$ in $M$ clusters by applying $M$-means clustering to $\{\tilde{{m}}_{1},\ldots,\tilde{{m}}_{\tilde{M}}\}$ using Lloyd’s Algorithm
@@
    2 ${{{\pi}}^{(j)}}=\sum_{i\in{\mathcal{I}}_{j}}{\tilde{{{\pi}}}^{{(i)}}}$; ${m}_{j}=\frac{1}{{{{\pi}}^{(j)}}}\sum_{i\in{\mathcal{I}}_{j}}{\tilde{{{\pi}}}^{{(i)}}}\tilde{{m}}_{i}$; $\Sigma_{j}=\frac{1}{{{{\pi}}^{(j)}}}\sum_{i\in{\mathcal{I}}_{j}}{\tilde{{{\pi}}}^{{(i)}}}\left[\tilde{\Sigma}_{i}+(\tilde{{m}}_{i}-{m}_{j})(\tilde{{m}}_{i}-{m}_{j})^{T}\right]$
 3Return $q=\sum_{j=1}^{M}{{{\pi}}^{{(j)}}}\mathcal{N}({m}_{j},\Sigma_{j})$
@@
Compression Operation
Our approach to compress a GMM of size $N$ into a GMM of size $M<N$, i.e., operation ${G}_{{M_{k}}}$ in Eqn (19b), is described in Algorithm 3 and is based on the M-means clustering using Lloyd’s algorithm (Lloyd, 1982) on the means of the mixture’s components (line 1). That is, each cluster is substituted by a Gaussian distribution with mean and covariance equal to those of the mixture in the cluster (line 2).
Remark 25
For the compression of discrete distributions, such as in the case of Bayesian inference via dropout, we present a computationally more efficient alternative procedure to Algorithm 3 in Section E of the Appendix. Note that in the dropout case, the support of the multivariate Bernoulli distribution representing the dropout masks at each layer’s input grows exponentially with the layer width, making compression of paramount importance in practice.
@@ 7.2 Construction of a SNN Approximation of a GMM
The algorithmic framework presented in the previous subsection finds the Gaussian Mixture distribution that best approximates the output distribution of a SNN and returns bounds on their distance. In this subsection, we show how our results can be employed to solve the complementary problem of finding the parameter distribution of a SNN to best match a given GMM.
Let us consider a finite set of input points ${\mathcal{X}}\subset{\mathbb{R}}^{n_{0}}$ and a SNN and GMM, whose distributions at $\mathcal{X}$} are respectively $p_{nn(\mathcal{X})}$ and $q_{gmm({\mathcal{X}})}$. Furthermore, in this subsection, we assume that the parameter distribution $p_{{w}}$ of the SNN is parametrized by a set of parameters $\psi$. For instance, in the case where $p_{{w}}$ is a multivariate Gaussian distribution, then $\psi$ parametrizes its mean and covariance. Note that, consequently, also $p_{nn({\mathcal{X}})}$ depends on $\psi$ through Eqn (24). To make the dependence explicit, in what follows, we will use a superscript to emphasize the dependency on $\psi$. Then, our goal is to find $\psi^{*}=\mathrm{argmin}_{\psi}{\mathbb{W}}_{2}(p_{nn(\mathcal{X})}^{\psi},q_{gmm(\mathcal{X})})$. To do so, we rely on Corollary 26 below, which uses the triangle inequality to extend Corollary 20.
Corollary 26 (of Corollary 20)
@@
${\mathbb{W}}_{2}\big{(}p_{nn({\mathcal{X}})}^{\psi},q_{gmm({\mathcal{X}})}\big{)}\leq\hat{{\mathbb{W}}}_{2}\big{(}p_{nn({\mathcal{X}})}^{\psi},q_{nn({\mathcal{X}})}^{\psi}\big{)}+{\mathbb{MW}}_{2}\big{(}q_{nn({\mathcal{X}})}^{\psi},q_{gmm({\mathcal{X}})}\big{)},$ (32)
where the GMM $q_{nn({\mathcal{X}})}^{\psi}$ and the bound $\hat{{\mathbb{W}}}_{2}$ are obtained via Algorithm 1, and ${\mathbb{MW}}_{2}$ is as defined in Definition 2, with $q_{nn({\mathcal{X}})}^{\psi}$ also depending on $\psi$ through Eqn (25).
@@
$\psi^{*}\approx\underset{\psi}{\mathrm{argmin}}\big{\{}\beta\hat{{\mathbb{W}}}_{2}\big{(}p^{\psi}_{nn({\mathcal{X}})},q^{\psi}_{nn({\mathcal{X}})}\big{)}+(1-\beta){\mathbb{MW}}_{2}\big{(}q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\big{)}\big{\}},$ (33)
where $\beta\in[0,1]$ allows us to trade off between the gap between the bound $\hat{{\mathbb{W}}}_{2}\big{(}p^{\psi}_{nn({\mathcal{X}})},q^{\psi}_{nn({\mathcal{X}})}\big{)}$ and ${\mathbb{W}}_{2}\big{(}p^{\psi}_{nn({\mathcal{X}})},q^{\psi}_{nn({\mathcal{X}})}\big{)}$ and the gap between the bound ${\mathbb{MW}}_{2}\big{(}q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\big{)}$ and ${\mathbb{W}}_{2}\big{(}q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\big{)}$. For instance, in the case where $q^{\psi}_{nn({\mathcal{X}})}$ and $q_{gmm({\mathcal{X}})}$ are Gaussian distributions, ${\mathbb{MW}}_{2}\big{(}q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\big{)}$ equals ${\mathbb{W}}_{2}\big{(}q^{\psi}_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\big{)}$, leading us to choose $\beta<\frac{1}{2}$. As discussed next, under the assumption that the mean and covariance of $p_{{w}}$ are differentiable with respect to $\psi$, the objective in Eqn (33) is piecewise differentiable with respect to $\psi$. Hence, the optimization problem can be approximately solved using gradient-based optimization techniques, such as Adam (Kingma and Ba, 2014).
Let us first consider the term $\hat{{\mathbb{W}}}_{2}$ in the objective, which is iteratively defined in Eqn (21) over the layers of the network using the quantities ${\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}]^{1/2}$, $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)$, and ${\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)$ for each layer $k$. In Algorithm 1, we compute these quantities so that the gradients w.r.t. the mean and covariance of $p_{{w}}$ are (approximately) analytically tractable:, as shown next:
@@
    ${\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}]^{1/2}$ is computed as described in Appendix C (line 6 of Algorithm 1) as a function of the sum of the variances of the $k$-th layer’s parameters and the spectral norm of the mean of the $k$-th layer’s weight matrix. Hence, the gradient with respect to the variance is well defined, and the gradient with respect to the mean exists if the largest singular value is unique; otherwise, we can use a sub-gradient derived from the singular value decomposition (SVD) of the mean matrix (Rockafellar, 2015).
@@
    $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)$ is obtained according to Algorithm 2 (line 4 of Algorithm 1) based on the eigendecomposition of the covariance matrix for each component of the GMM ${G}_{M_{k}}(q_{nn({\mathcal{X}}),k})$ (line 1 of Algorithm 2). Differentiability of $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)$ with respect to the mean and covariance of $p_{{w}}$ follows from the piecewise differentiability of the eigenvalue decomposition of a (semi-)positive definite matrix (Magnus and Neudecker, 2019), [^15] Note that there are discontinuous changes in the eigenvalues for small perturbations if the covariance matrix is (close to) degenerate, which occurs in our cases if the behavior of the SNN is strongly correlated for points in $\mathcal{X}$}. To prevent this, we use a stable derivation implementation for degenerate covariance matrices as provided by Kasim (2020) to compute the eigenvalue decomposition. provided that the covariances of ${G}_{M_{k}}(q_{nn({\mathcal{X}}),k})$ are differentiable with respect to the mean and covariance of $p_{{w}}$. To establish the latter, note that the compression operation ${G}_{{M_{k}}}$, as performed according to Algorithm 3, given the clustering of the components of the mixture $q_{nn({\mathcal{X}}),k}$, reduces to computing the weighted sum of the means and covariances of each cluster. Hence, the covariances of ${G}_{M_{k}}(q_{nn({\mathcal{X}}),k})$ are piecewise differentiable with respect to the means and covariances of $q_{nn({\mathcal{X}}),k}$. Here, $q_{nn({\mathcal{X}}),k}$ is an affine combination of $p_{{{w}}_{k-1}}$ and $\mathcal{X}$} if $k=1$, and the support of the signature approximation of ${G}_{M_{k}}(q_{nn({\mathcal{X}}),k-1})$ otherwise (see Example 1 for the closed-form expression of $q_{nn({\mathcal{X}}),k}$ in the case ${\mathcal{X}}=\{{{x}}\}$). As such, from the chain rule, it follows that for all $k$ the means and covariances of ${G}_{M_{k}}(q_{nn({\mathcal{X}}),k})$, and consequently, $\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right)$, are piecewise differentiable with respect to the means and covariances of $p_{{{w}}_{k-1}},\ldots,p_{{{w}}_{0}}$.